OpenStudy (anonymous):

Your gold team shot rounds of 68,70,72,78, and 84. If par for the course was 71, how many shots was the team over?

OpenStudy (yellowlegoguy99):

2xy = 10 - 3x y = (10 - 3x)/(2x)

OpenStudy (yellowlegoguy99):

17 over par.

OpenStudy (yellowlegoguy99):

Medal please?

OpenStudy (anonymous):

Ok. Thank you very much